Suppose that an estimate of Petco salaries is based on a sample of 27 employees. If we wish to conduct a test of whether the average salaries are over $35 thousand at an alpha of .01 , we would obtain the t-multiple using the Excel command:

a) = T.INV.2T(.02,27)
b) = T.INV.2T(.01,27)
c) = T.INV.2T(.01,26)
d) = T.INV.2T(.005,27)
e) None of the above
Answer #1

here for degree of freedom =n-1=27-1 =26

also as it is one tailed test ; probabiliy used must be 0.01*2 =0.02

therefore option E is correct (none of the above)